# Radiation-dominated polar emitting region of an accreting X-ray pulsar – I. Polarization- and spectrum-dependent structure, and the emergent continuum

### v1

@@ Abstract
The radiation-dominated polar emitting region of an accreting X-ray pulsar is simulated numerically in the framework of a three-dimensional (geometrically two-dimensional) model. The radiative transfer within the emitting region and the structure of the latter are calculated with the use of the self-consistent algorithm developed earlier. The magnetic scattering cross sections dependent on the photon energy and polarization have been incorporated. Second-order bulk Comptonization over entire emitting region, induced Compton scattering, the switching of the polarization modes, free-free processes, the cyclotron emission because of electron-proton collisions, and a realistic shape of the accretion channel have been taken into account. The case of a dipole magnetic field is considered. It is shown that the induced Compton effect can play a notable role in establishing the electron temperature in the post-shock zone. Within the model shock wave, a higher electron temperature is achieved than in the post-shock zone by means of the bulk-heating mechanism. The photons, gained the energy in the shock wave and above it due to bulk motion effects and the thermal Doppler effect, are responsible for the formation of high-energy regions in the emergent continuum of the polarization modes.
@@ 1 Introduction
An accreting X-ray pulsar (Giacconi et al., 1971) is a magnetized neutron star in a binary system on to which a plasma inflows from a companion star (e.g. Pringle & Rees 1972; Davidson & Ostriker 1973; Baan & Treves 1973). The accreting material is directed towards the neutron star along the magnetic field lines, increasing the kinetic energy, the main fraction of which is eventually released as the energy of X-ray photons emerging from heated to corresponding temperatures regions of the atmosphere of the neutron star or those grown above the star. Usually, the areas of the emitting regions across the magnetic field are much smaller than the area of the neutron star surface, and the magnetic and rotational axes of the neutron star are far enough apart to coincide. Hence, the rotation of the neutron star provides the possibility of the formation of the pulses of the X-ray radiation that comes from those regions to a distant observer.
@@
It is well known that modelling the deceleration of the accretion flow in a radiation-dominated shock wave is an example of a problem, which have not yet been considered exhaustively in the framework of a unitary study. The 2D numerical solution of Davidson (1973) became the first quantitative theoretical result showed the structure of the emitting region formed in sufficiently luminous accretion-powered X-ray pulsars. A one-dimentional approximation of the initially 2D equations from this work was considered by Basko & Sunyaev (1976). Blandford & Payne (1981a) derived in the Fokker-Planck approximation the radiative transfer equation (Chandrasekhar, 1960; Mihalas, 1978) for the compressed flow of a plasma. This result is used in most of the papers on the spectral radiative transfer in the accretion columns listed in the next two paragraphs. In the papers by Blandford & Payne (1981b) and Lyubarskii & Syunyaev (1982), the solutions for the spectrum of radiation of a 1D radiation-dominated shock were studied.
@@
Burnard, Arons & Klein (1991) calculated the spectra of the normal polarization modes from a preset radiation-dominated mound with taking into account the angular photon redistribution using the Feautrier method (Burnard, Klein & Arons, 1988, 1990), and the processes in the shock wave and above were neglected. The inclusion of the photon bubble oscillations in the numerical calculations of the radiation-dominated accretion column was performed by Klein & Arons (1989) (see also Arons 1992; Klein et al. 1996; Hsu, Arons & Klein 1997). A steady-state 1D dynamical model of the cylindrical column was developed by Becker (1998). Analytical solutions for the radiative transfer equation represented in the Fokker-Planck approximation for such a column were obtained by Becker & Wolff (2005, 2007) with making use of the velocity profiles allowing the separation of variables, but different from linear ones. Numerical solution of the spatially 1D radiative transfer equation with the use of the fenomenological power-law velocity profiles was carried out by Farinelli et al. (2012, 2016). Mushtukov et al. (2015a) obtained an expression for the critical luminosity and some solutions describing the structure of the column. West, Wolfram & Becker (2017a, b) self-consistently simulated the column taking into consideration the equations of state for the two-component plasma. The structure of the flow was supposed to be one-dimensional. Gornostaev (2021) (hereafter referred to as G2021) solved numerically the problems of the 3D dynamical structure of the inclined accretion columns, and the geometrically 2D problem of the structure of the column and its emergent spectrum in a self-consistent way, regardless of the straightforward observational applications (the gas pressure was neglected). The MHD simulations of the 2D structure of the column were performed by Zhang, Blaes & Jiang (2021, 2022), and development of the photon bubble instability was confirmed. Becker & Wolff (2022) took into account the upward widening of the accretion channel, analysed different velocity profiles (including those for the gas-dominated column), and added to their previous considerations some calculations of the cross sections in a strong magnetic field.
@@
It seems that a physically consistent modelling of the radiative transfer and accretion column structure taken into consideration the entire diversity of non-stationary processes, inhomogeneities in the gas flow, the influence of the gas component, and included the calculation of cyclotron processes and polarization of the radiation is a multidimensional problem which still does not have complete solution (Arons et al., 1987). Many years have passed since the estimates of Davidson & Ostriker (1973) appeared to be useful for thorough progress in understanding the nature of the observed phenomena. There are the observational features of the accretion-powered X-ray pulsars that need to be explained in the framework of theoretical models embracing as much properties of the real objects as possible. The formation of the spectrum within the polar emitting regions is related with two interesting manifestations (among others): variations in the continuum and in the parameters of the cyclotron resonant scattering feature (CRSF) (Truemper et al., 1978). It is remarkable that the experimental evidences obtained suggest the existence of the relationship between the changes in the continuum and CRSF (Klochkov et al., 2011).
@@
This is often associated the positive correlation of the CRSF energy with the X-ray luminosity $L_{\rm X}$ and the decrease (increase) in the height of the emitting region (being in the inhomogeneous magnetic field of an X-ray pulsar) with the increase (decrease) in $L_{\rm X}$ (Staubert et al., 2007; Rothschild et al., 2017; Lyutikov & Gavriil, 2006). Such a property of the emitting region is often explained (Rothschild et al., 2017; Vybornov et al., 2017; Staubert et al., 2019) by the formation of the gas-dominated accretion column bounded from above by a collisionless shock wave (Bisnovatyi-Kogan & Fridman 1970; Shapiro & Salpeter 1975; Langer & Rappaport 1982; de Hoffmann & Teller 1950). Mushtukov et al. (2015b) proposed the mechanism for the positive correlation related to the Doppler shift of the CRSF during transitions between regimes of a relatively rapid flow and the radiation-dominated accretion column. The negative correlation of the CRSF energy with $L_{\rm X}$ (Makishima et al., 1990; Tsygankov et al., 2006; Doroshenko et al., 2017; Vybornov et al., 2018) is usually related with the emitting region being in the radiation-dominated regime. However, there is no complete explanation for the phenomenon, and the number of the observed sources demonstrating such a dependence is very small. Possible known interpretations include: variations in the characteristic magnetic field strength due to an increase (decrease) in the height of the emitting region with increasing (decreasing) $L_{\rm X}$; variations in the magnetic field strength within the accretion funnel due to the MHD instabilities (Mukherjee, Bhattacharya & Mignone, 2013a); a significant influence on the CRSF position of the intra-atmospheric magnetic field of the neutron star during reflecting the direct radiation from the accretion funnel containing a column of variable height (Poutanen et al., 2013).
@@ 2.1 Basic equations
$+\left(\frac{kT}{\epsilon}\right)^{3}K_{i}{\rm e}^{-\frac{\epsilon}{kT}}\left(1-n_{i}\left({\rm e}^{\frac{\epsilon}{kT}}-1\right)\right)+j_{{\rm cyc,}\,i}=0,$
$i=1,~2,$
where $T$ is the electron temperature, $\hat{D}_{i}$ is the diffusion tensor, $\bar{\sigma}_{i}$ is the angle-averaged scattering cross section, $\sigma_{3-i\leftarrow i}$ is the cross section of the mode switching from the mode denoted by the right-hand subscript, $m_{\rm e}$ is the electron mass, $k$ is the Boltzmann constant, $K_{i}$ is the rate of free-free processes, and $j_{{\rm cyc},\,i}$ is the rate of change of the photon occupation number due to the production of cyclotron photons by means of de-excitations of the first Landau level because of electron-proton collisions. (The $i$ subscript takes everywhere below the values 1 and 2.) Apart from the mentioned processes of two latter lines of (5), these equations thus describe the spatial diffusion, advection, first-order bulk Comptonization (which attended by the compression of the plasma flow), thermal Comptonization, and second-order bulk Comptonization. The latter is described by adding the ${m_{\rm e}v^{2}}/{3}$ quantity to the electron temperature within the Doppler term in the Kompaneets operator (Psaltis & Lamb 1997; Titarchuk, Mastichiadis & Kylafis 1997; Farinelli et al. 2012, 2016). (As is known, bulk Comptonization of both kinds is associated with the Fermi processes of the non-thermal nature, ‘accelerating’ photons.)
@@
Ignoring the specific behaviour of the magnetic electron scattering cross sections near the cyclotron resonance, let us write them for the continuum below the cyclotron energy as (e.g. Canuto et al. 1971; Lodenquai et al. 1974; Arons et al. 1987)
@@
for the extraordinary mode and
$\sigma_{2}=\left(\sin^{2}\alpha+\left(\frac{\epsilon}{\epsilon_{\rm cyc}}\right)^{2}\cos^{2}\alpha\right)\sigma_{\rm T}$ (8)
for the ordinary mode, where $\epsilon_{\rm cyc}$ is the cyclotron energy at a given $r$, $\alpha$ is the angle between the wavevector of the incoming photon and the direction of the magnetic field, and $\sigma_{\rm T}$ is the Thomson cross section. For a medium dominated by the electron scattering (this is the case under consideration, as will be estimated and numerically verified), it is set
$\sigma_{\|,\,1}=\sigma_{\perp,\,1}=\bar{\sigma}_{1}=\left(\frac{\epsilon}{\epsilon_{\rm cyc}}\right)^{2}\sigma_{\rm T},$ (9)
$\sigma_{\|,\,2}=\left(\frac{\epsilon}{\epsilon_{\rm cyc}}\right)^{2}\sigma_{\rm T},~~\sigma_{\perp,\,2}=\sigma_{\rm T},$
@@
Simple estimates for a non-magnetized static medium show that the Compton scattering is a much more efficient mechanism of the energy exchange than free-free processes in the problem under consideration. Indeed, the rate of the Compton interactions is (Illarionov & Syunyaev, 1975a)
@@
In the computations, the delta function is approximated by the normalised Gaussian function
@@
where $\tilde{\varsigma}=\varsigma/1~{\rm keV}=1$ (e.g. Farinelli et al. 2016; West et al. 2017b). Normalisation is formally used for the infinite range of the variable, since the distribution is centred far from zero. Using normalisation for the semi-infinite range does not affect the results remarkably. It is assumed that $j_{\rm cyc,\,1}=j_{\rm cyc,\,2}$.
@@
or $v(r_{\rm up})$ is set to be equal to a certain value. As the upper boundary condition for the transfer equations, it is set that
@@ 2.2 Problem parameters
The mass and radius of the neutron star are $1.4M_{\odot}$ and $10^{6}$ cm, correspondingly. In what follows, the case of $l=3$ corresponding to the dipole magnetic field is considered. It is supposed that $T_{*}=1~{\rm keV}$, which is of the order of magnitude equal to the real electron temperature in the flow above the emitting region. It is set that $\gamma=0.98$. The $v(r_{\rm up})$ quantity is considered as a parameter in the particular case, where $v(r_{\rm up})=1.3\times 10^{10}$ cm s-1. Higher values of the mass accretion rate as well as distinct values of the surface magnetic field strength, the radius of the funnel base and the thickness of the filled part of the hollow funnel will be considered in a separate paper.
@@ 2.3 Numerical approach
The codes for simulating the accretion columns were written in C and executed in parallel with the use of OpenMP. The calculations were performed using total capabilities of two 56-core (AMD EPYC 7663) CPUs. The compute time of all simulations presented is about three months.
@@ 3.1 Accretion column in the radiation-dominated regime
The dynamical picture of the problem is known. Consider a freely falling flow. Let X-ray photons be generated at the base of the funnel, and let their total luminosity be determined by the expression (1) (a separate consideration of the case when photons are produced in the region of a certain height above the atmosphere is omitted). An increase in the accretion rate and the X-ray luminosity causes an increase of the role of the pressure of the radiation propagating upward. The flow begins to slow higher up, and the X-ray photons escaping the funnel begin to partially gain the energy in the appearing region of small values of $|\nabla\cdot\bm{v}|$, and even somewhat above it, in the near-divergence-free flow. A further increase in the accretion rate provides an increase in the fraction of photons gaining considerable energy in the shock wave and thus an increasingly sharp convergence of the flow. Once a sufficiently strong shock wave is formed, the settling mound begins to arise, as a significant portion of the photons penetrate beneath the shock. The seed free-free radiation is always originated mostly inside the underlying atmosphere. Photons appeared inside the settling mound accomplish random walks before escaping after a time equal on average to the mean diffusion time for the mound. As the accretion rate increases further, the shock wave becomes steeper and the peak height of the deceleration zone grows. Both effects occur (up to achieving some state), since the profile of the shock wave must provide the illumination of the radiation for a given accretion luminosity. A relatively small part of the pre-shock kinetic energy of protons is converted into the thermal energy of electrons. Within the shock, there is a surface where the supersonic decelerating flow becomes subsonic (the speed of sound in a radiation-dominated plasma is $c_{\rm s}=2/3\sqrt{u/(m_{\rm p}n_{\rm e})}$).
@@
$T_{\rm eff}=\left(\frac{\dot{M}v^{2}(r_{\rm up})}{2Ab}\right)^{\frac{1}{4}},$ (31)

### v3

@@ Abstract
The radiation-dominated polar emitting region of an accreting X-ray pulsar is simulated numerically in the framework of a three-dimensional (geometrically two-dimensional) model. The radiative transfer within the emitting region and the structure of the latter are calculated with the use of the self-consistent algorithm developed earlier. The magnetic scattering cross sections dependent on the photon energy and polarization have been incorporated. Second-order bulk Comptonization over entire emitting region, induced Compton scattering, the switching of the polarization modes, free-free processes, the cyclotron emission because of electron-proton collisions, and a realistic shape of the accretion channel have been taken into account. The case of a dipole magnetic field is considered. It is shown that the induced Compton effect can play a notable role in establishing the electron temperature in the post-shock zone. Within the model shock wave, a higher electron temperature is achieved than in the post-shock zone by means of the bulk-heating mechanism. The photons gaining the energy in the shock wave and above it due to bulk motion effects and the thermal Doppler effect are responsible for the formation of high-energy regions in the emergent continuum of the polarization modes.
@@ 1 Introduction
An accreting X-ray pulsar (Giacconi et al., 1971) is an included in a binary system, magnetized neutron star on to which a plasma inflows from a companion star (e.g. Pringle & Rees 1972; Davidson & Ostriker 1973; Baan & Treves 1973). The accreting material is directed towards the neutron star along the magnetic field lines, increasing the kinetic energy, the main fraction of which is eventually released as the energy of X-ray photons emerging from heated to corresponding temperatures regions of the atmosphere of the neutron star or those grown above the star. Usually, the areas of the emitting regions across the magnetic field are much smaller than the area of the neutron star surface, and the magnetic and rotational axes of the neutron star are far enough apart to coincide. Hence, the rotation of the neutron star provides the possibility of the formation of the pulses of the X-ray radiation that comes from those regions to a distant observer.
@@
It is well known that modelling the deceleration of the accretion flow in a radiation-dominated shock wave is an example of a problem, that has not yet been considered exhaustively in the framework of a unitary study. The 2D numerical solution of Davidson (1973) became the first quantitative theoretical result showed the structure of the emitting region formed in sufficiently luminous accretion-powered X-ray pulsars. A one-dimentional approximation of the initially 2D equations from this work was considered by Basko & Sunyaev (1976). Blandford & Payne (1981a) derived in the Fokker-Planck approximation the radiative transfer equation (Chandrasekhar, 1960; Mihalas, 1978) for the compressed flow of a plasma. This result is used in most of the papers on the spectral radiative transfer in the accretion columns listed in the next two paragraphs. In the papers by Blandford & Payne (1981b) and Lyubarskii & Syunyaev (1982), the solutions for the spectrum of radiation of a 1D radiation-dominated shock were studied.
@@
Burnard, Arons & Klein (1991) calculated the spectra of the normal polarization modes from a preset radiation-dominated mound with taking into account the angular photon redistribution using the Feautrier method (Burnard, Klein & Arons, 1988, 1990), and the processes in the shock wave and above were neglected. The inclusion of the photon bubble oscillations in the numerical calculations of the radiation-dominated accretion column was performed by Klein & Arons (1989) (see also Arons 1992; Klein et al. 1996; Hsu, Arons & Klein 1997). A steady-state 1D dynamical model of the cylindrical column was developed by Becker (1998). Analytical solutions for the radiative transfer equation represented in the Fokker-Planck approximation for such a column were obtained by Becker & Wolff (2005, 2007) with making use of the velocity profiles allowing the separation of variables, but different from linear ones. Numerical solution of the spatially 1D radiative transfer equation with the use of the fenomenological power-law velocity profiles was carried out by Farinelli et al. (2012, 2016). Mushtukov et al. (2015a) obtained an expression for the critical luminosity and some solutions describing the structure of the column. West, Wolfram & Becker (2017a, b) self-consistently simulated the column taking into consideration the equations of state for the two-component plasma. The structure of the flow was supposed to be one-dimensional. Gornostaev (2021) (hereafter referred to as G2021) solved numerically the problems of the 3D dynamical structure of the inclined accretion columns, and (in a self-consistent way) the geometrically 2D problem of the structure of the column and its emergent spectrum, regardless of the straightforward observational applications (the gas pressure was neglected). The MHD simulations of the 2D structure of the column were performed by Zhang, Blaes & Jiang (2021, 2022), and development of the photon bubble instability was confirmed. Becker & Wolff (2022) took into account the upward widening of the accretion channel, analysed different velocity profiles (including those for the gas-dominated column), and added to their previous considerations some calculations of the cross sections in a strong magnetic field.
@@
It seems that a physically consistent modelling of the radiative transfer and accretion column structure taking into consideration the entire diversity of non-stationary processes, inhomogeneities in the gas flow, the influence of the gas component, and including the calculation of cyclotron processes and polarization of the radiation is a multidimensional problem which still does not have complete solution (Arons et al., 1987). Many years have passed since the estimates of Davidson & Ostriker (1973) appeared to be useful for thorough progress in understanding the nature of the observed phenomena. There are the observational features of the accretion-powered X-ray pulsars that need to be explained in the framework of theoretical models embracing as much properties of the real objects as possible. The formation of the spectrum within the polar emitting regions is related with two interesting manifestations (among others): variations in the continuum and in the parameters of the cyclotron resonant scattering feature (CRSF) (Truemper et al., 1978). It is remarkable that the experimental evidences obtained suggest the existence of the relationship between the changes in the continuum and CRSF (Klochkov et al., 2011).
@@
The positive correlation of the CRSF energy with the X-ray luminosity $L_{\rm X}$ is often associated with the decrease (increase) in the height of the emitting region (being in the inhomogeneous magnetic field of an X-ray pulsar) at increasing (decreasing) $L_{\rm X}$ (Staubert et al., 2007; Rothschild et al., 2017; Lyutikov & Gavriil, 2006). Such a property of the emitting region is often explained (Rothschild et al., 2017; Vybornov et al., 2017; Staubert et al., 2019) by the formation of the gas-dominated accretion column bounded from above by a collisionless shock wave (Bisnovatyi-Kogan & Fridman 1970; Shapiro & Salpeter 1975; Langer & Rappaport 1982; de Hoffmann & Teller 1950). Mushtukov et al. (2015b) proposed the mechanism for the positive correlation related to the Doppler shift of the CRSF during transitions between regimes of a relatively rapid flow and the radiation-dominated accretion column. The negative correlation of the CRSF energy with $L_{\rm X}$ (Makishima et al., 1990; Tsygankov et al., 2006; Doroshenko et al., 2017; Vybornov et al., 2018) is usually related with the emitting region being in the radiation-dominated regime. However, there is no complete explanation for the phenomenon, and the number of the observed sources demonstrating such a dependence is very small. Possible known interpretations include: variations in the characteristic magnetic field strength due to an increase (decrease) in the height of the emitting region with increasing (decreasing) $L_{\rm X}$; variations in the magnetic field strength within the accretion funnel due to the MHD instabilities (Mukherjee, Bhattacharya & Mignone, 2013a); a significant influence on the CRSF position of the intra-atmospheric magnetic field of the neutron star during reflecting the direct radiation from the accretion funnel containing a column of variable height (Poutanen et al., 2013).
@@ 2.1 Basic equations
$+\left(\frac{kT}{\epsilon}\right)^{3}K_{i}{\rm e}^{-\frac{\epsilon}{kT}}\left(1-n_{i}\left({\rm e}^{\frac{\epsilon}{kT}}-1\right)\right)$
$+j_{{\rm cyc,}\,i}=0,~i=1,2,$
where $T$ is the electron temperature, $\hat{D}_{i}$ is the diffusion tensor, $\bar{\sigma}_{i}$ is the angle-averaged scattering cross section, $\sigma_{3-i\leftarrow i}$ is the cross section of the mode switching from the mode denoted by the right-hand subscript, $m_{\rm e}$ is the electron mass, $k$ is the Boltzmann constant, $K_{i}$ is the rate of free-free processes, and $j_{{\rm cyc},\,i}$ is the rate of production of cyclotron photons by means of de-excitations of the first Landau level because of electron-proton collisions. (The $i$ subscript takes everywhere below the values 1 and 2.) Apart from the mentioned processes of three latter lines of (5), these equations thus describe the spatial diffusion, advection, first-order bulk Comptonization (which attended by the compression of the plasma flow), thermal Comptonization, and second-order bulk Comptonization. The latter is described by adding the ${m_{\rm e}v^{2}}/{3}$ quantity to the electron temperature within the Doppler term in the Kompaneets operator (Psaltis & Lamb 1997; Titarchuk, Mastichiadis & Kylafis 1997; Farinelli et al. 2012, 2016). (As is known, bulk Comptonization of both kinds is associated with the Fermi processes of the non-thermal nature, ‘accelerating’ photons.)
@@
Ignoring the specific behaviour of the electron scattering cross sections near the cyclotron resonance, let us write them for the continuum below the cyclotron energy as (e.g. Canuto et al. 1971; Lodenquai et al. 1974; Arons et al. 1987)
@@
and
$\sigma_{2}=\left(\sin^{2}\alpha+\left(\frac{\epsilon}{\epsilon_{\rm cyc}}\right)^{2}\cos^{2}\alpha\right)\sigma_{\rm T},$ (8)
where $\epsilon_{\rm cyc}$ is the cyclotron energy at a given $r$, $\alpha$ is the angle between the wavevector of the incoming photon and the direction of the magnetic field, and $\sigma_{\rm T}$ is the Thomson cross section. For a medium dominated by the electron scattering (this is the case under consideration, as will be estimated and numerically verified), it is set
$\sigma_{\|,\,1}=\sigma_{\perp,\,1}=\bar{\sigma}_{1}=\sigma_{\|,\,2}=\left(\frac{\epsilon}{\epsilon_{\rm cyc}}\right)^{2}\sigma_{\rm T},~\sigma_{\perp,\,2}=\sigma_{\rm T},$ (9)
@@
Simple estimates for a non-magnetized static medium show that the Compton scattering is a much more efficient mechanism of the energy exchange than free-free processes in the problem under consideration. Indeed, the corresponding rate of the Compton interactions is (Illarionov & Syunyaev, 1975a)
@@
In the computations, the delta function is approximated by the normalized Gaussian function
@@
where $\tilde{\varsigma}=\varsigma/1~{\rm keV}=1$ (e.g. Farinelli et al. 2016; West et al. 2017b). Normalization is formally used for the infinite range of the variable, since the distribution is centred far from zero. Using normalization for the semi-infinite range does not affect the results remarkably. It is assumed that $j_{\rm cyc,\,1}=j_{\rm cyc,\,2}$.
@@
or $v(r_{\rm up})$ is set to be equal to a certain value. For the transfer equations, at the upper boundary it is set
@@ 2.2 Problem parameters
It is set that $M=1.4M_{\odot}$ and $R=10^{6}$ cm. In what follows, the case of $l=3$ corresponding to the dipole magnetic field is considered. It is supposed that $T_{*}=1~{\rm keV}$, which is of the order of magnitude equal to the real electron temperature in the flow above the emitting region. It is set that $\gamma=0.98$. The $v(r_{\rm up})$ quantity is considered as a parameter in the particular case, where $v(r_{\rm up})=1.3\times 10^{10}$ cm s-1.
Higher values of the mass accretion rate as well as distinct values of the surface magnetic field strength, the radius of the funnel base and the thickness of the filled part of the hollow funnel will be considered in a separate paper.
@@ 2.3 Numerical approach
The codes for simulating the accretion columns were written in C and executed in parallel with the use of OpenMP. The calculations were performed using total capabilities of two 56-core (AMD EPYC 7663) CPUs. The overall compute time of the presented simulations is about three months.
@@ 3.1 Accretion column in the radiation-dominated regime
The dynamical picture of the problem is known. Consider a freely falling flow. Let X-ray photons be generated at the base of the funnel, and let their total luminosity be determined by the expression (1) (a separate consideration of the case when photons are produced in the region of a certain height above the atmosphere is omitted). An increase in the accretion rate and the X-ray luminosity causes an increase of the role of the pressure of the radiation propagating upwards. The flow begins to slow higher up, and the X-ray photons escaping the funnel begin to partially gain the energy in the appearing region of small values of $|\nabla\cdot\bm{v}|$, and even somewhat above it, in the near-divergence-free flow. A further increase in the accretion rate provides an increase in the fraction of photons gaining considerable energy in the shock wave and thus an increasingly sharp convergence of the flow. Once a sufficiently strong shock wave is formed, the settling mound begins to arise, as a significant portion of the photons penetrate beneath the shock. The seed free-free radiation is always originated mostly inside the underlying atmosphere. Photons appeared inside the settling mound accomplish random walks before escaping after a time equal on average to the mean diffusion time for the mound. As the accretion rate increases further, the shock wave becomes steeper and the peak height of the deceleration zone grows. Both effects occur (up to achieving some state), since the profile of the shock wave must provide the illumination of the radiation for a given accretion luminosity. A relatively small part of the pre-shock kinetic energy of protons is converted into the thermal energy of electrons. Within the shock, there is a surface where the supersonic decelerating flow becomes subsonic (the speed of sound in a radiation-dominated plasma is $c_{\rm s}=2/3\sqrt{u/(m_{\rm p}n_{\rm e})}$).
@@
$T_{\rm eff}\approx\left(\frac{\dot{M}v^{2}(r_{\rm up})}{2Ab}\right)^{\frac{1}{4}},$ (31)